OpenStudy (anonymous):

Help!:)how to use summation formulas to rewrite equations without summation notation???

OpenStudy (anonymous):

\[\left[ 1-(1/4)^2 \right]+\left[ 1-(2/4)^2 \right]+...\left[ 1-(4/4)^2 \right]\]

OpenStudy (anonymous):

use sigma notation to write the sum?

OpenStudy (anonymous):

and use summation formulas to reqrite witout summation notaiton for \[\sum_{i=1}^{n} (4i+3)/n ^{2}\]

OpenStudy (anonymous):

for that one, you can tell the equation is 1-(i/4)^2 i figured out the i because that part was the only part that changed. so, you have to plug those into the sigma notation. it would be ∑ (with the 4 on top and i=1 on bottom) (1-(i/4)^2)

OpenStudy (anonymous):

for the second one, simplify the sigma notation. so you'd get something like 4/n^2∑i +1/n^2∑3. then, if you use the summation formulas, you should be able to simplify that so there'll be no ∑ signs in the answer

OpenStudy (anonymous):

cool thanks!:)